# A Classification Benchmark for Artificial Intelligence Detection of Laryngeal Cancer from Patient Speech

### v1

@@ 4.2 Fair Classification
In this section, we perform statistical tests comparing the correctly and incorrectly classified patients and their gender and age. First, a Fisher Exact test was used to compare the gender of correctly and incorrectly classified samples. This was done only on the FEMH holdout test set. Table 6 shows the p values from the Fisher Exact test for all classifiers. All of the p values are below the commonly used threshold of 0.05, meaning that there is a statistically significant relationship between the gender of the speaker and the performance of the classifier. This may be due to the imbalance of male and female participants across the malignant and benign groups. When investigating the patients that were correctly and incorrectly classified, we found male patients were misclassified more often than female patients across all models. This is likely because there are more male patients in both classes, whereas female patients are mainly in the benign class. Female patients with a malignant pathology were often misclassified as having a benign pathology due to this imbalance.
Model Input FeatureStates OpenSmile MFCC
SVM Voice 4.54e-33 2.44e-37 8.13e-22
Voice + Demographics 6.38e-40 1.17e-55 7.54e-40
Voice + Symptoms 1.48e-21 3.56e-22 1.36e-28
Voice + Demographics + Symptoms 5.72e-35 2.01e-15 1.06e-36
MLP Voice 2.60e-15 7.87e-52 1.10e-26
Voice + Demographics 3.26e-30 1.80e-56 7.56e-27
Voice + Symptoms 1.55e-15 1.08e-16 2.37e-19
Voice + Demographics + Symptoms 1.19e-11 2.79e-26 1.67e-21
Logistic Regression Voice 1.13e-22 1.50e-47 1.06e-13
Voice + Demographics 6.39e-25 6.72e-83 3.19e-56
Voice + Symptoms 3.13e-24 4.03e-18 1.45e-25
Voice + Demographics + Symptoms 4.51e-46 2.37e-37 5.52e-40
@@
Table 6: The p values from the Fisher Exact test comparing the male and female patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold.
@@
We performed a t-test to compare the ages of the correctly and incorrectly classified patients. The p-values are shown in Table 7. For all models, the p-value is below the standard 0.05 threshold, indicating a significant difference between the ages of the correctly and incorrectly classified patients. Further investigation found that younger patients are more often correctly classified than older patients. This is likely because young patients are almost exclusively found in the benign group and are therefore easier to classify.
Model Input FeatureStates OpenSmile MFCC
SVM Voice 1.75e-11 1.39e-05 6.15e-04
Voice + Demographics 3.70e-23 2.33e-28 2.40e-29
Voice + Symptoms 9.17e-10 6.89e-06 1.52e-06
Voice + Demographics + Symptoms 1.34e-20 8.40e-20 9.08e-13
MLP Voice 2.19e-07 7.57e-05 1.23e-04
Voice + Demographics 4.84e-14 7.02e-27 7.61e-21
Voice + Symptoms 6.45e-06 9.45e-05 1.74e-04
Voice + Demographics + Symptoms 3.43e-09 3.86e-13 7.52e-11
Logistic Regression Voice 2.97e-12 7.09e-04 1.77e-02
Voice + Demographics 2.79e-18 8.65e-33 3.82e-28
Voice + Symptoms 7.77e-10 2.11e-10 4.94e-07
Voice + Demographics + Symptoms 3.95e-31 5.60e-21 7.10e-13
@@
Table 7: The p values from the t-test comparing the age of the patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold.
@@
There is a statistical significance in the age and gender of correctly and incorrectly classified patients, even when voice alone is used as the system input. This suggests that there are audio features that may be indicative of a patient’s age and gender even when this information is not explicitly entered into the classifier. It is well known that the voices of males and females vary in pitch, although there can be a significant overlap (Latinus and Taylor, 2012). There is also an expected difference in voice as people age, with a change in pitch and the variation in pitch and amplitude (Rojas et al., 2020). These differences may cause some differences in classifier performance based on a patient’s sex or age without the information being specifically input into the system.
Graphs showing the difference in correctly and incorrectly classified male and female patients, as well as the age distribution of correctly and incorrectly classified patients, can be found in the GitHub repository and the appendix.
@@ 4.3 Prediction Time
In this section, we report the time taken to make a prediction for a single audio file for all of the presented models.
Figure 5 shows the time it takes to get a prediction from a single audio file, including feature extraction. The FeatureStates classifiers take significantly longer than the OpenSmile and MFCC classifiers due to the feature extraction. Overall, inference time does not vary significantly between the inputs or algorithms.
@@
Feature extraction takes the majority of the time taken to obtain a prediction. Figure 6 shows the inference times taken for samples after feature extraction. It is clear that even after feature extraction, the FeatureStates classifiers take the longest and MFCC the shortest amount of time, although the time taken for prediction is very short across all algorithms. The FeatureStates is the largest feature set, comprising 512 features. OpenSmile comprises 88, and MFCC comprises 20 coefficients but varies based on the recording length. When using the average length of an MFCC in the training set and flattening it, this comprises of 1540 features. The MFCCs are may be processed faster due to them being stored as numpy arrays rather than dataframes. However, since the difference is in the range of hundreds of seconds, this is unlikely to make a real-world difference.
We believe that since the total inference times, including feature extraction, are consistently less than 2 seconds, these inference times are short enough to be reasonably implemented into clinical practice.
Figure 6: The time required to make a prediction from an audio file after feature extraction. Outliers have been excluded from this figure.
@@ 5 Conclusion and Future Work
This work presents 36 models for classifying malignant and benign voice pathologies based on patients’ speech, demographics, and symptoms. We report the model’s performance, inference times, and statistical tests comparing classifier performance to patient demographics. All training and evaluation code, pre-trained models, and figure generation are available in a public repository for use as a baseline so that future work may be improved.
We find that while voice alone can be used to classify benign and malignant patients, the addition of demographic and symptom data improves classifier performance (69.1% balanced accuracy with voice only, 83.7% balanced accuracy with voice, demographics, and symptoms). We also find that while these classifiers may perform well on a holdout test set, performance degrades on external data.
In future work, we hope to improve these baseline models and their generalizability and robustness to improve their implementation in clinical practice. We hope to enhance the fairness of models in terms of patient’s age and gender. We hope that, as more datasets are release, these models may be further evaluated and enhanced.

### v2

@@ 4.2 Fairness
In this section, we perform statistical tests comparing the correctly and incorrectly classified patients and their sex and age. First, a Fisher Exact test was used to compare the sex of correctly and incorrectly classified samples. This was done only on the FEMH holdout test set. Table 9 shows the p values from the Fisher Exact test for all classifiers as well as the percentage of male and female patients that were misclassified. All of the p values are below the commonly used threshold of 0.05, meaning that there is a statistically significant relationship between the sex of the speaker and the performance of the classifier. This may be due to the imbalance of male and female participants across the malignant and benign groups. We can also see that male patients are misclassified more often than female patients across all models. This is likely because there are more male patients in both classes, whereas female patients are mainly in the benign class. Female patients with a malignant pathology were often misclassified as having a benign pathology due to this imbalance.
Inputs Algorithm Feature Fisher Exact P Value Female Patients Incorrectly Classified Male Patients Incorrectly Classified
Voice SVM FeatureStates 4.54e-33 11.8% 54.7%
OpenSmile 2.44e-37 9.4% 54.3%
MFCC 8.13e-22 4.2% 31.3%
MLP FeatureStates 2.60e-15 6.8% 30.2%
OpenSmile 7.87e-52 7.6% 61.2%
MFCC 1.10e-26 16.0% 55.4%
Logistic Regression FeatureStates 1.13e-22 11.8% 45.7%
OpenSmile 1.50e-47 9.7% 61.9%
MFCC 1.06e-13 16.2% 42.4%
Voice + Demographics SVM FeatureStates 6.38e-40 6.0% 50.0%
OpenSmile 1.17e-55 3.7% 55.8%
MFCC 7.54e-40 3.4% 44.2%
MLP FeatureStates 3.26e-30 4.7% 39.9%
OpenSmile 1.80e-56 3.1% 55.0%
MFCC 7.56e-27 2.4% 31.3%
Logistic Regression FeatureStates 6.39e-25 10.5% 45.7%
OpenSmile 6.72e-83 1.0% 64.7%
MFCC 3.19e-56 2.4% 52.9%
Voice + Symptoms SVM FeatureStates 1.48e-21 5.5% 33.8%
OpenSmile 3.56e-22 5.2% 33.8%
MFCC 1.36e-28 9.2% 46.8%
MLP FeatureStates 1.55e-15 6.5% 29.9%
OpenSmile 1.08e-16 7.3% 32.4%
MFCC 2.37e-19 10.2% 39.9%
Logistic Regression FeatureStates 3.13e-24 8.4% 41.7%
OpenSmile 4.03e-18 7.6% 34.5%
MFCC 1.45e-25 8.1% 42.4%
Voice + Demographics + Symptoms SVM FeatureStates 5.72e-35 4.2% 42.4%
OpenSmile 2.01e-15 11.3% 37.4%
MFCC 1.06e-36 7.1% 49.6%
MLP FeatureStates 1.19e-11 1.3% 14.7%
OpenSmile 2.79e-26 4.2% 35.3%
MFCC 1.67e-21 2.6% 27.3%
Logistic Regression FeatureStates 4.51e-46 2.1% 45.3%
OpenSmile 2.37e-37 1.6% 37.4%
MFCC 5.52e-40 4.5% 46.8%
@@
Table 9: A table showing the p values from the Fisher Exact test comparing the male and female patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold. This table also shows the percentage of male and female patients that were incorrectly classified by each model. The green highlighted rows indicate the best performing models for each combination of input variables.
@@
We then performed a t-test to compare the ages of correctly and incorrectly classified patients. The corresponding p-values are shown in Table 10 as well as the mean age of the correctly and incorrectly classified patients. For all models, the p-value is below the standard 0.05 threshold, indicating a significant difference between the ages of correctly and incorrectly classified patients. It can be seen that younger patients are more often correctly classified than older patients. This is likely because young patients are almost exclusively found in the benign group and are therefore easier to classify.
Inputs Algorithm Feature T Test P Value Average Age of Incorrect Classification Average Age of Correct Classification
Voice SVM FeatureStates 1.75e-11 55 46
OpenSmile 1.39e-05 53 47
MFCC 6.15e-04 53 48
MLP FeatureStates 2.19e-07 55 47
OpenSmile 7.57e-05 52 47
MFCC 1.23e-04 52 47
Logistic Regression FeatureStates 2.97e-12 56 46
OpenSmile 7.09e-04 52 47
MFCC 1.77e-02 51 48
Voice + Demographics SVM FeatureStates 3.70e-23 59 45
OpenSmile 2.33e-28 59 45
MFCC 2.40e-29 61 45
MLP FeatureStates 4.84e-14 58 46
OpenSmile 7.02e-27 59 45
MFCC 7.61e-21 62 46
Logistic Regression FeatureStates 2.79e-18 57 46
OpenSmile 8.65e-33 59 44
MFCC 3.82e-28 60 45
Voice + Symptoms SVM FeatureStates 9.17e-10 56 47
OpenSmile 6.89e-06 54 47
MFCC 1.52e-06 54 47
MLP FeatureStates 6.45e-06 55 47
OpenSmile 9.45e-05 54 48
MFCC 1.74e-04 53 47
Logistic Regression FeatureStates 7.77e-10 55 47
OpenSmile 2.11e-10 56 47
MFCC 4.94e-07 54 47
Voice + Demographics + Symptoms SVM FeatureStates 1.34e-20 59 46
OpenSmile 8.40e-20 58 46
MFCC 9.08e-13 56 46
MLP FeatureStates 3.43e-09 61 48
OpenSmile 3.86e-13 58 47
MFCC 7.52e-11 59 47
Logistic Regression FeatureStates 3.95e-31 62 45
OpenSmile 5.60e-21 61 46
MFCC 7.10e-13 56 46
@@
Table 10: The p values from the t-test comparing the age of the patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold. The mean age of the correctly and incorrectly classified patients is also shown. The green highlighted rows indicate the best performing models for each combination of input variables.
@@
There is a statistical significance in the age and sex of correctly and incorrectly classified patients, even when the demographic data is not being used as input to the model. This suggests that there are audio features that may be indicative of a patient’s age and sex even when this information is not explicitly entered into the classifier. It is well known that the voices of males and females vary in pitch, although there can be significant overlap Latinus and Taylor (2012). There is also an expected difference in voice as people age, with a change in pitch and the variation in pitch and amplitude Rojas et al. (2020). These differences in voice may cause unfair classifier performance based on a patient’s sex or age without this data being directly input into the system.
For the problem of classifying benign and malignant patients, fairness is an issue. Creating fair classifiers is difficult especially due to the known lack of female laryngeal cancer patients Cancer Research UK (2021b). This demographic disparity means that creating classifiers that are able to accurately classify female patients is especially difficult. The average age of laryngeal cancer patients is also high meaning that young patients are also likely to be under classified as having malignant pathologies. There is significant work to be done to develop classifiers that are fairer.
@@ 4.3 Prediction Time
There has been a recent shift towards the use of large and complex AI systems. While these systems sometimes achieve impressive results, it’s important to understand how they would be practically implemented into practice. In a clinical setting it’s essential that these systems do not slow down referral processes. As such we feel that it’s vital that prediction times are reported. This is currently rarely done in this area of research.
In this section, we report the time taken to make a prediction for a single audio file for all of the presented models in order to understand whether these models would cause delays in patient referrals. Figure 6 shows the time it takes to get a prediction from a single audio file for each model. It can be seen that the FeatureStates classifiers take significantly longer than the OpenSMILE and MFCC classifiers. The inference time does not vary significantly between the inputs or algorithms.
@@
Further investigation found that the feature extraction stage takes most of the time to obtain a prediction for all models. This explains why the FeatureStates models take significantly longer than the OpenSMILE and MFCC models to make predictions since the extraction of features from a large neural network takes much longer than extracting features using OpenSMILE or converting audio to MFCCs.
All models take less than two seconds to make a prediction for a single audio file. We believe that the total inference times are short enough to be reasonably implemented into clinical practice without slowing the patient referral process.
@@ 5 Conclusion and Future Work
This work presents a benchmark suite of 36 models for classifying malignant and benign voice pathologies based on patients’ voice, demographics, and symptoms. The classification pipeline used in this work can be used for future work with each element changed to optimize model performance. We also suggest standardized model evaluation techniques to allow the easier comparison of results within this research area.
For the 36 presented models we report the model’s predictive performance, statistical fairness tests, and inference times. All training and evaluation code, trained models, and figure generation are available in a public repository for use as a baseline so that future work may be improved. All models are trained and tested on publicly available datasets meaning that these results can be directly reproduced. The availability of code allows for further evaluation on future datasets and for these models to be used as a baseline for future researchers to work from.
The results of our models show that while voice alone can be used to classify benign and malignant patients, the addition of demographic and symptom data improves classifier performance (69.1% balanced accuracy with voice only, 83.7% balanced accuracy with voice, demographics, and symptoms). We also evaluate the performance of these models on an external test set and find that while there is some loss in performance, it is not significant showing that these models are generalizable.
When comparing these results to other work using the same datasets to perform the same classification task we find that our methods outperform more complex deep learning methods. The most significant performance improvement is that of the model using audio features only achieving an increase in balanced accuracy of 14.6% on the holdout test set and 17.3% on the external test set. These results show that these simpler machine learning algorithms when combined with feature extraction can outperform deep learning methods.
We also evaluated the practicality of these models in terms of clinical implementation. First we used statistical tests to evaluate the fairness of these models in regards to patient’s age and sex. We found that these models are not fair likely due to the significant disparity in the demographics of patients in the benign and malignant groups. In future work this must be addressed to ensure that systems are able to be implemented with confidence that they are fair. Next we reported the inference times of each model. We found that all models can make predictions in less than two seconds, which we feel would not slow down patient referral.
In future work, we hope to improve the predictive performance of these baseline models as well as improving generalizability and robustness. We hope to enhance the fairness of models in terms of patient’s age and sex especially as this is an area that the presented models are lacking. We hope that, as more datasets are release, these models may be further evaluated and enhanced. We also hope that in future work the robustness of these models may be further evaluated in terms of the presence of background noise and other variations to the recording environment so that we can be sure that models could be successfully transitioned into clinical practice. We also hope that the technique used to develop these models may be translated beyond the detection of laryngeal cancer into the detection of other vocal pathologies such as predicting specific benign pathologies such as vocal cord palsies, vocal nodules, and benign vocal cord lesions.
Acknowledgements
This research was funded in part by the UKRI Engineering and Physical Sciences Research Council (EPSRC) [EP/S024336/1].